OpenStudy (anonymous):

Expand the logarithm expression. Log6 b/8

OpenStudy (jack1):

hay @amandaleonhardt the best page i would recommend for understnding this is http://www.mathsisfun.com/algebra/logarithms.html

OpenStudy (jack1):

now ur qustion... is it supposed to be : \[\large \log 6 = b/8 ?\] or \[\large \log_6 \frac b8 ?\]